# Myers–Steenrod theorems for metric and singular Riemannian foliations

### v1

@@ 3. Proof of Theorem A
Let $p\colon X\to Y$ be a proper submetry from a length space $X$ to a metric space $Y$ and set =$\mathcal{F}=\{p^{-1}(y)\mid y\in Y\}$. If $\operatorname{Isom}(X)$ admits a Lie group structure with respect to the open-compact topology and $\operatorname{Isom}(Y)$ is closed, then $\operatorname{Isom}(X,\mathcal{F})$ is a Lie group.
3.1. Proof of Theorem A (i)
Let $(M,\mathcal{F})$ be a singular Riemannian foliation with closed leaves on a complete Riemannian manifold and equip the leaf space $M/\mathcal{F}$ with the leaf-projection metric. As recalled in Section 2.3, the leaf-projection map $\pi\colon M\to M/\mathcal{F}$ is a submetry.
@@ 3.1. Proof of Theorem A (i)
The leaf-projection map $\pi\colon M\to M/\mathcal{F}$ is proper.
@@
Let $K$ be a compact subset of the leaf space $M/\mathcal{F}$. Since $M/\mathcal{F}$ is a complete metric space, $K$ is totally bounded. Thus, given $\varepsilon>0$, we may cover $K$ with a finite number of closed balls of radius $\varepsilon$. By the slice theorem for singular Riemannian foliations (see [34, Theorem A]), we can choose $\varepsilon$ in such a way that each $\varepsilon$-ball in the cover corresponds to a closed tubular neighborhood of some leaf. Thus, the preimage of $K$ is contained in a finite union of compact sets, and is therefore compact. Hence, the leaf-projection map is proper. ∎
@@
The group of isometries of $M/\mathcal{F}$ is a closed subgroup of $\operatorname{\mathrm{Homeo}}(M/\mathcal{F})$.
@@
Let $\{f_{n}^{\ast}\}\subset\operatorname{Isom}(M/\mathcal{F})$ be a sequence with limit $f^{\ast}\in\operatorname{\mathrm{Homeo}}(M/\mathcal{F})$. Let us show that $f^{\ast}\in\operatorname{Isom}(M/\mathcal{F})$. Consider two distinct points $p^{\ast}$ and $q^{\ast}$ in $M/\mathcal{F}$. Applying twice the triangle inequality, we obtain
$d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))\leq d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f_{n}^{\ast}(p^{\ast}))+d_{M/\mathcal{F}}(f_{n}^{\ast}(p^{\ast}),f_{n}^{\ast}(q^{\ast}))+d_{M/\mathcal{F}}(f_{n}^{\ast}(q^{\ast}),f^{\ast}(q^{\ast})).$
Taking $\varepsilon>0$, assuming that $n$ is sufficiently large, and using the fact that each $f^{\ast}_{n}$ is an isometry of $M/\mathcal{F}$, we obtain that
$d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))$ $\leq d_{M/\mathcal{F}}(f_{n}^{\ast}(p^{\ast}),f_{n}^{\ast}(q^{\ast}))+2\varepsilon$
$=d_{M/\mathcal{F}}(p^{\ast},q^{\ast})+2\varepsilon.$
Thus,
(3.1) $d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))-d_{M/\mathcal{F}}(p^{\ast},q^{\ast})\leq 2\varepsilon.$
Similarly, using again the fact that $f_{n}^{*}$ is an isometry of $M/\mathcal{F}$ and applying twice the triangle inequality, we obtain
$d_{M/\mathcal{F}}(p^{\ast},q^{\ast})\leq d_{M/\mathcal{F}}(f_{n}^{\ast}(p^{\ast}),f^{\ast}(p^{\ast}))+d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))+d_{M/\mathcal{F}}(f^{\ast}(q^{\ast}),f_{n}^{\ast}(q^{\ast})).$
Assuming again that $n$ is sufficiently large, we obtain that
$d_{M/\mathcal{F}}(p^{\ast},q^{\ast})\leq d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))+2\varepsilon.$
Thus,
(3.2) $d_{M/\mathcal{F}}(p^{\ast},q^{\ast})-d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))\leq 2\varepsilon.$
Inequalities (3.1) and (3.2) imply that, for any $\varepsilon>0$,
$|d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))-d_{M/\mathcal{F}}(p^{\ast},q^{\ast})|\leq 2\varepsilon.$
This implies that $d_{M/\mathcal{F}}(f^{\ast}(p^{\ast}),f^{\ast}(q^{\ast}))=d_{M/\mathcal{F}}(p^{\ast},q^{\ast})$. Since $p^{\ast}$ and $q^{\ast}$ are arbitrary, we conclude that $f^{\ast}$ is an isometry. ∎
With the preceding observations in hand, we may now prove the first item in Theorem A.
Proof of Theorem A (i).
First, recall that the Riemannian metric of $M$ induces a metric $d^{\ast}$ on $M/\mathcal{F}$, making it a complete length space $(M/\mathcal{F},d^{\ast})$. Moreover, $(M/\mathcal{F},d^{\ast})$ is locally an Alexandrov space, as pointed out in Remark 2.19 (see [30, pp. 77–78]). Thus, by Lemma 3.3, the map $\pi\colon M\to M/\mathcal{F}$ is a proper submetry, and by Lemma 3.4, the group $\operatorname{Isom}(M/\mathcal{F})$ is closed. Moreover, by Theorem 2.18 (1), since $(M/\mathcal{F})$ is locally an Alexandrov space, $\operatorname{Isom}(M/\mathcal{F})$ is a Lie group. The desired conclusion now follows from Corollary 3.2. ∎
3.2. Proof of Theorem A (ii)
We first recall the following useful fact, which we will use in the proof of the theorem.
@@ 3.2. Proof of Theorem A (ii)
Let $f,g\colon X\to Y$ be continuous functions between topological spaces. If $Y$ is Hausdorff and there exists a dense subset $D\subset X$ of $X$ such that $f|_{D}=g|_{D}$, then $f=g$.

### v3

@@ 3. $(3)$
Let $p\colon X\to Y$ be a continuous surjective map between metric spaces and set $\mathcal{F}=\{p^{-1}(y)\mid y\in Y\}$}. If $\operatorname{Isom}(X)$ admits a Lie group structure with respect to the compact-open topology, then $\operatorname{Isom}(X,\mathcal{F})$ is a Lie group.
3 Proof of Theorem A
We will prove each item in Theorem A separately.
3.1 Proof of Theorem A(i)(i)
By the classical Myers–Steenrod theorem, $\operatorname{Isom}(M)$ is a Lie group and is compact if $M$ is compact. The conclusion now follows from Corollary 2.24.
3.2 Proof of Theorem A(ii)(ii)
Recall that $M_{\mathrm{prin}}\subset M$, the set of points in principal leaves, is an open and dense subset of $M$ (see Section 2.1). Hence, $M_{\mathrm{prin}}$ is a Riemannian manifold foliated by the principal leaves of $\mathcal{F}$, and the quotient $M_{\mathrm{prin}}/\mathcal{F}$ is also a Riemannian manifold. Moreover, the leaf-projection map $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$} is surjective and continuous. By the classical Myers–Steenrod theorem, $\operatorname{Isom}(M_{\mathrm{prin}})$ is a Lie group with respect to the compact-open topology. Hence, by Corollary 2.24, $\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$ is a Lie group.
Consider the group homomorphism
$\mathrm{P}\colon\ \operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$
given by $\mathrm{P}(f)=f|_{M_{\mathrm{prin}}}$. Note that $\mathrm{P}(f)\in\operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})$, since foliated isometries send principal leaves to principal leaves. Furthermore, $\mathrm{P}$ is continuous and, therefore, smooth by [31, Corollary 3.50]. Hence, $\mathrm{P}$ is a Lie group homomorphism.
Since $M_{\mathrm{prin}}/\mathcal{F}$ is a Riemannian manifold, $\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$ is a Lie group, by the classical Myers–Steenrod theorem. Consider the group homomorphism
$\Psi_{\mathrm{prin}}\colon\ \operatorname{Isom}(M_{\mathrm{prin}},\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F}),$
where $\Psi_{\mathrm{prin}}(f)$ is given by $\Psi_{\mathrm{prin}}(f)(\pi(x))=\pi(f(x))$ (cf. Section 2.7). The leaf-projection map $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$} is a fiber bundle with (nk)(n-k)-dimensional fibers (see [44, Theorem A] and [13, p. 41]). By Theorem 2.22, $\Psi_{\mathrm{prin}}$ is continuous and, therefore, smooth. Hence $\Psi_{\mathrm{prin}}$ is a Lie group homomorphism.
Consider now the Lie group homomorphism
$\Phi=\Psi_{\mathrm{prin}}\circ\mathrm{P}\colon\ \operatorname{Isom}(M,\mathcal{F})\to\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F}).$
By the first isomorphism theorem for Lie groups,
$\dim(\operatorname{Isom}(M,\mathcal{F}))=\dim(\ker(\Phi))+\dim(\Phi(\operatorname{Isom}(M,\mathcal{F})).$ (3.1)
Since $\Phi(\operatorname{Isom}(M,\mathcal{F}))$ is a Lie subgroup of $\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F})$ and $M_{\mathrm{prin}}/\mathcal{F}$ is a $k$-dimensional Riemannian manifold, by [33, Chapter II, Theorem 3.1],
$\dim(\Phi(\operatorname{Isom}(M,\mathcal{F})))\leq\dim(\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F}))\leq\frac{k(k+1)}{2}.$ (3.2)
Let $G=\ker(\Phi)\leq\operatorname{Isom}(M,\mathcal{F})$ and observe that $g\in G$ if and only if $g(L)=L$ for any principal leaf $L$. Fix $p\in M_{\mathrm{prin}}$ and write $T_{p}M=V_{p}\oplus H_{p}$, where $V_{p}=T_{p}L$ is the vertical space and $H_{p}=V_{p}^{\perp}$ is the horizontal space. For any $g\in G$, we have $\pi\circ g=\pi$, which implies that ${\rm d}\pi_{p}\circ{\rm d}g_{p}={\rm d}\pi_{p}$}. Since $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$ is a Riemannian submersion, ${\rm d}\pi_{p}|_{H_{p}}$ is a linear isomorphism onto $T_{\pi(p)}(M_{\mathrm{prin}}/\mathcal{F}$), and therefore ${\rm d}g_{p}|_{H_{p}}=\mathrm{Id}$. Hence, $G_{p}$, the isotropy group at $p$, acts trivially on $H_{p}$ and orthogonally on $V_{p}$. Therefore,
$\dim(G_{p})\leq\dim(\mathrm{O}(V_{p}))\leq\frac{(n-k)(n-k-1)}{2}.$ (3.3)
Since the orbit $G(p)$ of $G$ through $p$ is contained in the principal leaf containing $p$,
$\dim(G(p))\leq n-k.$ (3.4)
Since $\dim(G)=\dim(G(p))+\dim(G_{p})$, inequalities (3.3) and (3.4) imply that
$\dim(\ker(\Phi))=\dim(G)\leq n-k+\frac{(n-k)(n-k-1)}{2}=\frac{(n-k)(n-k+1)}{2}.$ (3.5)
Combining equation (3.1) with inequalities (3.2) and (3.5), we obtain
$\dim(\operatorname{Isom}(M,\mathcal{F}))\leq\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2},$ (3.6)
thus verifying inequality (a).
3.3 Proof of Theorem A(iii)(iii)
Assume equality holds in inequality (a), i.e.,
$\dim(\operatorname{Isom}(M,\mathcal{F}))=\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$
Then, by equation (3.1) and inequalities (3.2) and (3.5), we must have
$\dim(\Phi(\operatorname{Isom}(M,\mathcal{F})))=\frac{k(k+1)}{2},$ (3.7)
and
$\dim(\ker(\Phi))=\frac{(n-k)(n-k+1)}{2}.$ (3.8)
Inequality (3.2) and equation (3.7) imply that $\dim(\operatorname{Isom}(M_{\mathrm{prin}}/\mathcal{F}))=k(k+1)/2$. Since $M_{\mathrm{prin}}/\mathcal{F}$ is a $k$-dimensional Riemannian manifold, by [33, Chapter II, Theorem 3.1], $M_{\mathrm{prin}}/\mathcal{F}$ is isometric to one of the following $k$-dimensional space forms: Euclidean space, hyperbolic space, a round sphere, or round real projective space. In particular, $M_{\mathrm{prin}}/\mathcal{F}$ is complete. Since $M/\mathcal{F}$ is the metric completion of $M_{\mathrm{prin}}/\mathcal{F}$, it follows that $M/\mathcal{F}=M_{\mathrm{prin}}/\mathcal{F}$. Therefore, the foliation is regular and all the leaves are principal.
Since $\ker(\Phi)$ acts effectively by isometries on any leaf and leaves are (nk)(n-k)-dimensional, equation (3.8) implies that the isometry group of any leaf has maximal possible dimension. Hence, by [33, Chapter II, Theorem 3.1], that all the leaves are isometric to a round $S^{n-k}$, a round $\mathbb{R}P^{n-k}$, Euclidean space $\mathbb{R}^{n-k}$, or a hyperbolic space $\mathbb{H}^{n-k}$.
Recall that the leaf-projection map $\pi\colon M_{\mathrm{prin}}\to M_{\mathrm{prin}}/\mathcal{F}$ is a smooth Riemannian submersion. Since all the leaves are principal, $\pi\colon M\to M/\mathcal{F}$ is a fiber bundle, and thus the desired conclusion follows.
4 Proof of Theorem B
We will prove each item in Theorem B separately. Throughout this section, we let $\pi\colon X\to Y$ be a submetry between Alexandrov spaces and set $\mathcal{F}=\big\{\pi^{-1}(y)\mid y\in Y\big\}$. In the proofs of items $(ii)$ and $(iii)$ we assume $\pi$ is proper.
4.1 Proof of Theorem B(i)(i)
By Theorem 2.18(1)(1), $\operatorname{Isom}(X)$ is a Lie group and is compact if $X$ is compact. The conclusion now follows from Corollary 2.24.
4.2 Proof of Theorem B(ii)(ii)
We now compute the upper bound on the dimension of the Lie group $\operatorname{Isom}(X,\mathcal{F})$. Recall that from now on and in the rest of the section we assume that the submetry $\pi\colon X\to Y$ is proper. Suppose that $X$ has dimension $n\geq 1$ and $Y$ has dimension $0\leq k\leq n$, so that the regular leaves of the foliation $\mathcal{F}$ have dimension nkn-k. We will show that
$\dim(\operatorname{Isom}(X,\mathcal{F}))\leq\frac{k(k+1)}{2}+\frac{(n-k)(n-k+1)}{2}.$
Recall that a point $y\in Y$ is regular with respect to $\pi$ if $\pi$ is regular at each $x\in\pi^{-1}(y)$ (see [36, Definition 6.5]). The set of regular points
$Y_{\mathrm{reg}}=\{y\in Y\mid y\mbox{ is regular with respect to }\pi\}$
contains the open and dense subset of thick points (see [36, end of Section 6]). Consequently, the set $X_{\mathrm{reg}}=\{x\in X\mid x\mbox{ is regular}\}$ contains an open and dense subset. Let
$R_{X}=\big\{x\in X\mid\Sigma_{x}X\mbox{ is isometric to the unit round }S^{n-1}\big\}.$
By Theorem 2.12, $R_{X}$ is dense. Hence, $R_{X}\cap X_{\mathrm{reg}}\neq\varnothing$.
Fix $\bar{x}\in R_{X}\cap X_{\mathrm{reg}}$. We may choose $\bar{x}$ so that $\Sigma_{\pi(\bar{x})}Y$ is isometric to the unit round $S^{k-1}$, since the set of points in $Y$ whose space of directions is isometric to the unit round $S^{k-1}$ is dense in $Y$ and $Y_{\mathrm{reg}}$ contains an open subset. Since $\bar{x}$ is regular,
${\rm d}_{\bar{x}}\pi|_{H_{\bar{x}}}\colon\ H_{\bar{x}}\to\Sigma_{\pi(\bar{x})}Y$
is an isometry (see Section 2.4). Moreover, since $\Sigma_{\pi(\bar{x})}Y$ is round, there exists an isometry between $\Sigma_{\bar{x}}X$ and the spherical join $H_{\bar{x}}*V_{\bar{x}}$. Since $\Sigma_{\bar{x}}X$ is isometric to the unit round $S^{n-1}$ and $H_{\bar{x}}$ is isometric to the unit round $S^{k-1}$, we conclude that $V_{\bar{x}}$ is isometric to the unit round $S^{n-k-1}$.
Since $Y$ is an Alexandrov space, $\operatorname{Isom}(Y)$ is a Lie group, by Theorem 2.18. Consider the group homomorphism $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$, where $\Psi(f)$ is given by $\Psi(f)(\pi(x))=\pi(f(x))$ (cf. Section 2.7). We assume $\pi\colon X\to Y$ is proper. Hence, Theorem 2.22 implies that $\Psi$ is continuous, and thus it is a Lie group homomorphism.
Set $G=\ker(\Psi)$. Given $f\in G$, we have $f(\bar{x})\in\pi^{-1}(\pi(\bar{x}))$, allowing us to identify the differential ${\rm d}_{\bar{x}}f\colon\Sigma_{\bar{x}}X\to\Sigma_{f(\bar{x})}X$ (which is an isometry) with an isometry
$({\rm d}_{\bar{x}}f)_{1}*({\rm d}_{\bar{x}}f)_{2}\colon\ H_{\bar{x}}*V_{\bar{x}}\to H_{f(\bar{x})}*V_{f(\bar{x})}.$
Since $G$ acts on $\pi^{-1}(\pi(\bar{x}))$, we may consider the isotropy group $G_{\bar{x}}$. For $f\in G_{\bar{x}}$, we have then that $({\rm d}_{\bar{x}}f)_{1}=\mathrm{Id}_{H_{\bar{x}}}$ and $({\rm d}_{\bar{x}}f)_{2}\colon V_{\bar{x}}\to V_{\bar{x}}$ is an isometry. Thus, $G_{\bar{x}}$ acts by isometries on the unit round $S^{n-k-1}$. Moreover, this action is effective. Indeed, if $({\rm d}_{\bar{x}}f)_{2}=\mathrm{Id}_{V_{\bar{x}}}$, then ${\rm d}_{\bar{x}}f=\mathrm{Id}_{\Sigma_{\bar{x}}X}$. Since $f(\bar{x})=\bar{x}$ and $f$ is an isometry of the Alexandrov space $X$, then [23, Lemma 3.2] implies that $f=\mathrm{Id}_{X}$. Hence, by the Myers–Steenrod theorem,
$\dim(G_{\bar{x}})\leq\frac{(n-k-1)(n-k)}{2}.$
Now we consider the orbit $G(\bar{x})\subset\pi^{-1}(\pi(\bar{x}))$. By [36, Korollar 7.5], the Hausdorff dimension of $\pi^{-1}(\pi(\bar{x}))$ is nkn-k. Thus, we conclude that the dimension of $G(\bar{x})=G/G_{\bar{x}}$ is bounded above by nkn-k. It follows that
$\dim(G)=\dim(G(\bar{x}))+\dim(G_{\bar{x}})\leq n-k+\frac{(n\hskip-0.5pt-\hskip-0.5ptk\hskip-0.5pt-\hskip-0.5pt1)(n\hskip-0.5pt-\hskip-0.5ptk)}{2}=\frac{(n\hskip-0.5pt-\hskip-0.5ptk)(n\hskip-0.5pt-\hskip-0.5ptk\hskip-0.5pt+\hskip-0.5pt1)}{2}.\!\!\!\!$ (4.1)
$\dim(\Psi(\operatorname{Isom}(X,\mathcal{F})))\leq k(k+1)/2.$ (4.2)
Then, by the first isomorphism theorem for Lie groups, inequalities (4.1) and (4.2), and recalling that $G=\ker(\Psi)$,
$\dim(\operatorname{Isom}(X,\mathcal{F}))=\dim(G)+\dim(\Psi(\operatorname{Isom}(X,\mathcal{F})))\leq\frac{(n\hskip-0.5pt-\hskip-0.5ptk)(n\hskip-0.5pt-\hskip-0.5ptk\hskip-0.5pt+\hskip-0.5pt1)}{2}+\frac{k(k\hskip-0.5pt+\hskip-0.5pt1)}{2}.\!\!\!$ (4.3)
Remark 4.1.
The conclusions of Theorem A(ii)(ii) do not follow directly from the conclusions of Theorem B(ii)(ii), since the leaf space of a closed singular Riemannian foliation $\mathcal{F}$ on a complete manifold $M$ is only locally an Alexandrov space of bounded curvature (see [38, p. 119]). This is because the local lower curvature bound on $M/\mathcal{F}$ is given by a local lower sectional curvature bound on $M$. When $M$ is not compact, it may happen that $M$ does not have a global lower sectional curvature bound, and thus in this case $M/\mathcal{F}$ does not have a lower curvature bound in the sense of Alexandrov.
Remark 4.2.
Due to the lack of a slice theorem for submetries between Alexandrov spaces, we need properness of $\pi\colon X\to Y$ to guarantee that $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ is continuous, and thus a Lie group homeomorphism, to be able to say that $\dim(\operatorname{Isom}(X,\mathcal{F}))=\dim(\ker\Psi)+\dim(\operatorname{Image}(\Psi))$.
4.3 Proof of Theorem B(iii)(iii)
Recall our standing assumption that $\pi\colon X\to Y$ is a proper submetry between Alexandrov spaces. Assume that equality holds in (4.3), i.e.,
where $\dim(X)=n$ and $\dim(Y)=k$ with $0\leq k\leq n$. We first determine the topology of the base space $Y$.
@@ 4.3 Proof of Theorem B(iii)(iii)
The base space $Y$ is isometric to $k$-dimensional Euclidean space, hyperbolic space with constant negative sectional curvature, a round real projective space, or a round sphere.
@@
Since $Y$ is $k$-dimensional, Theorem 2.18 implies that $\dim(\operatorname{Isom}(Y))\leq k(k+1)/2$. Since $\dim(\operatorname{Isom}(X,\mathcal{F}))$ is maximal, it follows from the proof of Theorem B(ii)(ii) that the subgroup $\Psi(\operatorname{Isom}(X,\mathcal{F}))\subset\operatorname{Isom}(Y)$} has dimension $k(k+1)/2$. Thus, $\operatorname{Isom}(Y)$ has dimension $k(k+1)/2$ and the conclusion follows from Theorem 2.18. ∎
We now determine the topology of the fibers. Consider the Lie group homomorphism $\Psi\colon\operatorname{Isom}(X,\mathcal{F})\to\operatorname{Isom}(Y)$ with $\Psi(f)$ given by $\Psi(f)(\pi(x))=\pi(f(x))$ and set $G=\ker(\Psi)\subset\operatorname{Isom}(X,\mathcal{F})$.
@@
For any $x\in X$, the orbit $G(x)\subset\pi^{-1}(\pi(x))$ is a compact subset.
@@
Fix $x\in X$ and let $d_{I}$ be the intrinsic metric on $L_{x}=\pi^{-1}(\pi(x))$ induced by $d|_{L_{x}\times L_{x}}$. We first prove that $G\subset\operatorname{Isom}(L_{x},d_{I})$. Let $\gamma\colon[0,1]\to L_{x}$ be a rectifiable curve. Since $G\subset\operatorname{Isom}(X)$, for any $g\in G$, $L(g\circ\gamma)=L(\gamma)<\infty$, where $L(\cdot)$ denotes length. By definition,
$d_{I}(x,y)=\inf\{L(\gamma)\mid\gamma\colon[0,1]\to L_{x}\text{ is rectifiable with }\gamma(0)=x,\,\gamma(1)=y\}.$
Thus, $d_{I}(g(x),g(y))\leq d_{I}(x,y)$. Since $g$ is arbitrary,
$d_{I}(x,y)=d_{I}\bigl(g^{-1}(g(x)),g^{-1}(g(y))\bigr)\leq d_{I}(g(x),g(y)).$
Thus, $G\subset\operatorname{Isom}(L_{x},d_{I})$.
By construction, $G(x)\subset L_{x}$. We now show that $G(x)$ is closed. Consider $\{g_{i}\}_{i\in\mathbb{N}}\subset G$ with $\lim_{i\to\infty}g_{i}(x)=y\in X$}. By Proposition 2.21, the action of $\operatorname{Isom}(X)$ on $X$ is proper. Then, there exists a subsequence $\{g_{i_{k}}\}_{k\in\mathbb{N}}$ converging to some $g\in\operatorname{Isom}(X)$ (see [35, Proposition 21.5], whose proof is purely topological, and cf. [15, Proposition 2.2] or [18]). Recall that the map $\Psi$ is continuous. Hence, $G=\ker(\Psi)$ is closed in $\operatorname{Isom}(X,\mathcal{F})$, and $\operatorname{Isom}(X,\mathcal{F})$ is closed in $\operatorname{Isom}(X)$. Therefore, $g\in G$. By the continuity of the action of $G$ on $X$, $y=g(x)\in G(x)$. Thus, the $G$-orbits are closed in $X$. Since $\pi\colon X\to Y$ is proper, $L_{x}$ is compact, and since $G(x)\subset L_{x}$ is closed, $G(x)$ is also compact. ∎
Connected fiber case
We first consider the case where $\pi$ has connected fibers.
@@
If all fibers of $\pi\colon X\to Y$ are connected, then ((with the intrinsic metric)) each fiber is isometric to a round sphere or a round real projective space.